what is the formula for electro negativity in kilo joules/mole in paulings scale

Grade:12th pass

1 Answers

jitender
6 years ago
Xp=0.336 (Xm-0.615)Xm=(IP+EA)/2Or Xp=electronegativity on pauling scaleXm=on mullikan scaleIP=ionisation potentialEP=electron affinityPut ip and ep in kj/mol